OpenStudy (anonymous):

Use Cauchy's Theorem or Cauchy's Integral Theorem (or both) to evaluate the following. How does your answer change if C was to be a closed path composed of either or both straight line segments or circular arcs. HELP PLEASE?!

OpenStudy (anonymous):

\[\int\limits_{C}^{}\pi*\exp(\pi*z ^{-})dz\] where the integral is one with a circle on it, and z- is z-bar (z with a line above it)

OpenStudy (anonymous):

good question